OpenStudy (lgbasallote):

LGBADERIVATIVES!! \[\Large y = 2^{\sin \pi x}\]

OpenStudy (lgbasallote):

im thinking ln first

OpenStudy (konradzuse):

so this is the in form a^u

OpenStudy (konradzuse):

(ln|a|)a^u * u'

OpenStudy (lgbasallote):

\[\ln y = \ln (\sin \pi x) \ln 2\]

OpenStudy (helder_edwin):

\[ \large y'=2^{\sin\pi x}(\ln 2)(\cos\pi x)\pi \]

OpenStudy (ash2326):

We know that \[\frac{d}{dx} a^x=a^x \times \ln a\] Here we have \[\large y=2^{\sin (\pi x)}\] Let's differentiate this \[\frac{dy}{dx}= 2^{\sin (\pi x)} \times \ln 2 \times \frac{d}{dx} (\sin \pi x)\times \frac{d}{dx} (\pi x)\]

OpenStudy (lgbasallote):

\[\LARGE \frac 1y y' = \frac{1}{\sin \pi x} \cos \pi x \ln 2?\]

OpenStudy (helder_edwin):

there is no need for implicit derivatives

OpenStudy (lgbasallote):

no?

OpenStudy (helder_edwin):

just apply the chain rule!

OpenStudy (lgbasallote):

actually a^u is implicit derivatives....

OpenStudy (lgbasallote):

however i think there should be a 1/sin pi x somewhere...

OpenStudy (lgbasallote):

oh wait i did it wrong

OpenStudy (lgbasallote):

nevermind

OpenStudy (konradzuse):

i'm pretty sure a^u is supposed to be int here a swell.